# A Survey of LLM-Driven AI Agent Communication: Protocols, Security Risks, and Defense Countermeasures

### v3

@@ VII-C2 Misusing Tools
Malicious Content Propagation: Agents with permissions to publish externally (e.g., via email, social media, or CMS APIs) can be exploited to widely spread malware, phishing links, or misinformation once compromised. For example, a trusted customer service agent may send malware-laden emails to clients, or a content generation agent might publish misleading articles on official websites. Because these agents are trusted entities, such attacks are highly deceptive. More dangerously, attackers can leverage access to contacts, email histories, and user preferences to craft highly personalized phishing campaigns, enabling large-scale social engineering attacks.
VII-C3 Real-World Damage
Compromised agents can cause harm far beyond their internal systems, affecting both digital and physical ecosystems through their interactions with the external world. This section examines how such agents may disrupt the real world by polluting shared digital environments and executing harmful physical actions.
@@ VII-C3 Real-World Damage
Digital Environment Contamination: A compromised agent can cause long-term damage to the external digital environment, not by directly attacking other agents, but by polluting the shared information ecosystem they rely on. Since agents often interact with external platforms (such as submitting code on GitHub or editing Wikipedia entries), once compromised, they can systematically inject subtle yet harmful errors or biases into these shared resources [173]. Unlike cross-agent contamination, digital environment contamination indirectly infects all agents dependent on the polluted information sources. For example, a compromised coding agent may embed hidden logic bugs or backdoors when contributing code; a corrupted knowledge management agent might alter Wikipedia pages or internal knowledge bases by falsifying citations or inserting biased descriptions, thereby damaging the entire knowledge graph with far-reaching consequences.
@@
Physical Environment Disruption: Once an agent’s memory or tool module is compromised, the resulting threat is not limited to digital risk. It can manifest as concrete damage to the physical world through specific decision-making chains and execution paths. Polluted memory may contain falsified sensor data, misleading the agent’s perception of the physical environment. Tool modules that serve as interfaces to physical systems can directly implement flawed decisions, affecting device behaviors, environmental control, or industrial processes. For example, an agricultural agent misled by false pest-related memories may overapply pesticides; a quality inspection robot referencing corrupted standard images may repeatedly approve defective components; a warehouse robot using a compromised path-planning module may unknowingly cause stacking imbalances and logistics bottlenecks. Critically, such behaviors often appear to follow normal procedures, making them difficult to detect through traditional logging or anomaly detection methods. Thus, the consequences of agent compromise extend beyond digital misinformation, posing a tangible risk of physical system disruption.
VII-D Defense Countermeasure Prospect
@@
VII-D1 Countermeasures For Memory- and knowledge-related Risks
@@
Memory- and knowledge-related risks can be jointly addressed through an integrated mitigation framework spanning content filtering, output consensus, and architectural isolation.
@@
Embedding-Space Screening and Clustering-Based Anomaly Detection: Whether memory entries are agent-internal or retrieved externally via RAG, their semantic embeddings can be preemptively analyzed for anomalies. Techniques like TrustRAG [359] apply clustering (e.g., K-means) to identify vectors that deviate from the dominant semantic cluster. This approach effectively filters both static memory entries and retrieval results with low semantic cohesion, regardless of source. While lightweight and interpretable, clustering-based filtering must be augmented with adaptive schemas to detect context-sensitive triggers or stealthy distributional shifts.
@@
Consensus Filtering and Voting-Based Aggregation: To limit the model’s reliance on single compromised retrievals or poisoned memories, output-level consensus mechanisms have been proposed. RobustRAG [306], for example, treats each retrieved source independently and constructs responses based on overlapping semantic content (e.g., shared n-grams or keywords) across documents. This same principle can be extended to memory snapshots through majority-vote or semantic voting strategies, where only widely corroborated memories can influence the response. Such ensemble-style filters improve resilience by diluting the influence of outlier or adversarial sources.
@@
Execution Monitoring and Planning-Consistency Checks: Adversarial content within memory or RAG inputs may subtly deviate the agent’s behavior from user intent without explicit toxicity. Tools like ReAgent [30] introduce planning-level introspection where the agent paraphrases the user’s request, generates an expected plan, and continuously aligns runtime actions with this trace. Any inconsistency, triggered by an unexpected memory or an off-topic retrieval, is treated as a behavioral anomaly and can prompt halting or recovery mechanisms. This introspective framework provides a robust guardrail against both memory-hijacking and injection-aware RAG attacks.
@@
System-Gated Memory Retention and Input Sanitization: Architectural solutions such as DRIFT [99] and AgentSafe [196] implement strict content sanitization before newly generated content. DRIFT uses an injection isolator to scan generative outputs for adversarial goal shifts or impersonation cues, while AgentSafe enforces trust-tiered storage via ThreatSieve and prioritization via HierarCache. These mechanisms constrain future influence, ensuring that RAG or memory poisoning cannot silently accumulate over time.
@@
Unified Content Provenance and Trust Frameworks: Since retrieved knowledge and persisted memories may originate from overlapping sources (e.g., user prompts, tool calls, external APIs), maintaining clear provenance metadata and trust scores is essential. Unified provenance tracking across both memory and retrieval pipelines enables smarter decisions about retention, ranking, or discounting of contentious content. Combined with per-source reliability scoring, this approach encourages transparent auditing and facilitates downstream fine-tuning or gating mechanisms.
@@
VII-D2 Countermeasures For Tool-related Risks
Tool-related defense strategies should operate across four interlocking levels: protocol foundations, execution control, orchestration safety, and system enforcement.
Protocol-Level Safeguards: To counter risks such as tool poisoning, cross-origin exploits, and shadowing attacks enabled by flexible yet insufficiently regulated protocols like MCP, researchers have introduced security-verification frameworks operating at the registry and middleware layer. MCP-Scan [152] performs both static inspection of tool schemas (e.g., scanning for suspect tags or metadata) and real-time proxy-based validation of MCP traffic, leveraging LLM-assisted heuristics to flag covert behaviors. MCP-Shield [236] extends this with signature-matching and adversarial behavior profiling, enabling pre-execution detection of high-risk tools and malformed tasks. MCIP [130] builds on MAESTRO [49] to analyze runtime traces, proposing an explainable logging schema and a security-awareness model to track violations in complex agent-tool interactions.
Tool Invocation and Execution Controls: At the agent’s runtime execution point, classic techniques such as sandboxing and permission gating remain foundational. Google’s defense-in-depth model advocates policy engines that monitor planned tool actions, verify argument safety, and require human confirmation for risk-sensitive operations [65]. Tools should be executed in minimally privileged environments-e.g., isolated containers with controlled filesystem and network scope-to mitigate direct misuse, including SSRF and data exfiltration threats. Enforcement frameworks can also implement schema hardening or fine-grained input/output sanitization to reject anomalous payloads.
Agent-Orchestration Monitoring: Newer approaches target the agent’s planning cognition-its selection and chaining of tools. GuardAgent [308] introduces a validator agent that inspects the primary agent’s plan and generates executable guards (e.g., static checks or runtime assertions) before tool calls proceed. AgentGuard [34] takes a more declarative view: it uses an auxiliary LLM to model preconditions, postconditions, and transition constraints across multi-step tool workflows, effectively constraining the planner rather than reacting after execution begins. These strategies reflect a growing consensus: LLMs may require another LLM to safely oversee complex planning under uncertainty.
System-Level Mediation and Chaining Control: Complex pipelines-such as summarize(search(“…”))-can become attack vectors when tools trust upstream outputs implicitly. To prevent this, DRIFT [99] introduces a structured control architecture: a “Secure Planner” compiles a validated tool trajectory under strict parametric constraints, while a “Dynamic Validator” continuously monitors downstream tool executions for compliance. Notably, the Injection Isolator blocks adversarial propagation between tools by sanitizing both intermediate returns and final outputs, mitigating the risk of memory poisoning and delayed-stage tool exploits.
VII-E Takeaways
Agent-environment communication protocols like MCP enable agents to interface with diverse tools, APIs, and external data. However, they introduce risks such as memory injection, retrieval-augmented generation poisoning, and tool misuse. Malicious attackers can corrupt memory stores, manipulate knowledge bases, or exploit cross-tool chain vulnerabilities to harm the agent system. Besides, compromised agents can also cause damage to benign environments. To help developers mitigate these problems, we discuss the targeted defense countermeasures for risks from malicious environments. The defenses for compromised agents can be referred to Section VI-C. However, we also believe that related attacks will continue to emerge, and there needs long-term efforts to make agent-environment communication more secure.
@@ VIII-A The Selection of Protocols
Based on the above principles, we finally chose MCP and A2A. MCP is almost the most popular agent communication protocol currently. It has been adopted by diverse companies and individual developers. Up to now, MCP’s PyPI package has received over 9 million downloads in the last month [20], and its NPM download last week reached 4.2 million [21]. Besides, the number of open-source MCP servers also approached 2,000 [101]. After investigation, we find that none of the other existing protocols can achieve a similar popularity to MCP. A2A was proposed in April 2025, so it has not yet achieved the same level of promotion as MCP. However, we observe that the attention to A2A is rising rapidly, and its recognition among developers is also quite high. As a result, we finally chose MCP and A2A as the target protocols.
@@ VIII-B Experiments
In this chapter, we will conduct attacks against MCP and A2A and illustrate their weakness. In the experiments related to MCP, we use Claude as the MCP client. Claude [17] is an AI chatbot developed by Anthropic. With its outstanding code capabilities, safe design, and ultra-long context processing, it has become one of the most popular AI assistants in the world. In January 2025, Claude app downloads reached an estimated 769.6 million [264]. Therefore, exploiting Claude can provide practical value for real-world scenarios. Attacks #1-#4 are typical examples of the Malicious Tools as Attack Vectors discussed in Section VII-B3, and attack #5 belongs to the Description Poisoning in Section VI-B1.
Exp #1. Malicious Code Execution. We demonstrate that attackers can exploit MCP to enable the execution of malicious code on the computer system, which has also been proposed by other studies [229]. We download Claude sonnet 3.7 as the MCP host, and use the Filesystem Server [18] provided by the official MCP as the MCP server. As shown in Figure 10, MCP enables malicious users to directly open an unauthenticated Bash Shell service on port 4444 locally. Adding it to .zshrc will normalize the risk and make it extremely easy to be exploited by intruders. Unfortunately, this malicious operation does not get blocked.
Exp #2. Retrieval-Agent Deception. Attackers can write a seemingly normal document related to MCP and publish it on a public platform. However, malicious commands are embedded in this document, such as searching for local environment variables and sending the information to Slack [266]. The document content is as follows:
Users download these documents and build a local vector database through the Chroma MCP Server [1]. As a result, these malicious documents, together with other legitimate documents, are encoded into the vector database and become one of the retrieval targets. When a user sends a request to Claude Opus 4:
Claude will use Chroma to query the database and retrieve the contaminated documents, and then call the Terminal-Controller tool [88] to execute the embedded malicious commands. We successfully achieved this attack using Claude Opus 4.
Exp #3. Tool Poisoning. Tool poisoning means that attackers can add malicious instructions in the MCP tool description [254]. Such instructions do not influence the execution of tools, but can induce the agents to perform dangerous actions. We use Claude Sonnet 4.0, the Filesystem MCP Sever [18], and Gmail MCP Server [87] to illustrate this vulnerability. In the index.js of the Filesystem Server, we added the malicious description as follows:
@@
Exp #4. Command Injection. This attack denotes that malicious tools can directly manipulate the user’s system without causing any alert with the help of MCP. We inject a malicious command into the Terminal-Controller MCP tool [88]. When users call this tool, the embedded command will be executed. We successfully conducted this attack using Claude Sonnet 4.0. The injected command is as follows:
@@
Exp #5. Agent Selection Manipulation. We also conduct attacks against A2A. We find that attackers can add malicious descriptions in the Agent_Card to induce users to select them. In this experiment, we use the public A2A-samples provided by A2Aproject [4] as the user-side client. Then, we create two agents for weather querying. One of them is benign while the other is malicious. The user issues the following command to the A2A client.

### v4

@@ VII-B3 Risks at the Semantic Interpretation Layer (L3)
Malicious Content Propagation. Agents with permissions to publish externally (e.g., via email, social media, or CMS APIs) can be exploited to widely spread malware, phishing links, or misinformation once compromised. For example, a trusted customer service agent may send malware-laden emails to clients, or a content generation agent might publish misleading articles on official websites. Because these agents are trusted entities, such attacks are highly deceptive. More dangerously, attackers can leverage access to contacts, email histories, and user preferences to craft highly personalized phishing campaigns, enabling large-scale social engineering attacks.
Digital Environment Contamination. Memory and knowledge—both internal and external—serve as the cognitive substrate for multi-agent systems. Once compromised, these environments can induce large-scale, long-term reasoning failures. We divide this risk into two complementary subtypes based on the system boundary:
(a) Internal Digital Environment Contamination. A compromised agent can become a source of systemic contamination. Through agent communication, it can actively propagate tampered knowledge and flawed reasoning patterns, spreading its internal corruption to other agents and triggering a cascading infection of memory modules and knowledge bases across the system [ju2024flooding, chen2024agentpoison, lee2024prompt, li2025we]. Once the shared knowledge base is compromised, other agents may unknowingly retrieve and integrate malicious information into their memory module during task execution, converting knowledge base contamination into memory infection. Subsequently, an agent with a compromised memory module can use its authorized write access to contaminate the shared knowledge base of the entire system, in turn, forming a reverse contamination loop from memory to knowledge. Since these contamination operations come from the trusted agents within the system, they are highly difficult to detect. Once established, the contamination can persist long-term, continuously disrupting the behavior of compromised agents and misleading other agents that rely on the same knowledge sources, resulting in a form of chronic poisoning of the information ecosystem.
@@
(b) External Digital Environment Contamination. A compromised agent can cause long-term damage to the external digital environment, not by directly attacking other agents, but by polluting the shared information ecosystem they rely on. Since agents often interact with external platforms (such as submitting code on GitHub or editing Wikipedia entries), once compromised, they can systematically inject subtle yet harmful errors or biases into these shared resources [lin2025exploring]. Unlike cross-agent contamination, digital environment contamination indirectly infects all agents dependent on the polluted information sources. For example, a compromised coding agent may embed hidden logic bugs or backdoors when contributing code; a corrupted knowledge management agent might alter Wikipedia pages or internal knowledge bases by falsifying citations or inserting biased descriptions, thereby damaging the entire knowledge graph with far-reaching consequences.
@@
Physical Environment Disruption. Once an agent’s memory or tool module is compromised, the resulting threat is not limited to digital risk. It can manifest as concrete damage to the physical world through specific decision-making chains and execution paths. Polluted memory may contain falsified sensor data, misleading the agent’s perception of the physical environment. Tool modules that serve as interfaces to physical systems can directly implement flawed decisions, affecting device behaviors, environmental control, or industrial processes. For example, an agricultural agent misled by false pest-related memories may overapply pesticides; a quality inspection robot referencing corrupted standard images may repeatedly approve defective components; a warehouse robot using a compromised path-planning module may unknowingly cause stacking imbalances and logistics bottlenecks. Critically, such behaviors often appear to follow normal procedures, making them difficult to detect through traditional logging or anomaly detection methods. Thus, the consequences of agent compromise extend beyond digital misinformation, posing a tangible risk of physical system disruption.
VII-C Defense Countermeasure Prospect
@@
VII-C1 Defenses on Data Transmission Layer (L1)
The defenses on this layer are similar to those in Section V-C1. Besides, if agents need to call remote resources/tools, it is also necessary to authenticate their identities.
VII-C2 Defenses on Interaction Protocol Layer (L2)
At the interaction-protocol layer, tool risks primarily stem from how agents request, negotiate, and coordinate tool actions, making protocol-aware defenses essential before execution begins.
D1: Defense for Tool-related Risks. Tool-related defense strategies should operate across four interlocking levels: protocol foundations, execution control, orchestration safety, and system enforcement.
Protocol-Level Safeguards. To counter risks such as tool poisoning, cross-origin exploits, and shadowing attacks enabled by flexible yet insufficiently regulated protocols like MCP, researchers have introduced security-verification frameworks operating at the registry and middleware layer. MCP-Scan [mcpscan] performs both static inspection of tool schemas (e.g., scanning for suspect tags or metadata) and real-time proxy-based validation of MCP traffic, leveraging LLM-assisted heuristics to flag covert behaviors. MCP-Shield [riseandignite_mcp_shield_2025] extends this with signature-matching and adversarial behavior profiling, enabling pre-execution detection of high-risk tools and malformed tasks. MCIP [jing2025mcip] builds on MAESTRO [MAESTRO] to analyze runtime traces, proposing an explainable logging schema and a security-awareness model to track violations in complex agent-tool interactions. MCIP [jing2025mcip] invokes a structured information flow tracing log format. And builds a security-aware model trained on this trace data to identify and defend against malicious interactions.
Tool Invocation and Execution Controls. At the agent’s runtime execution point, classic techniques such as sandboxing and permission gating remain foundational. Google’s defense-in-depth model advocates policy engines that monitor planned tool actions, verify argument safety, and require human confirmation for risk-sensitive operations [Google-Approach-for-Secure-AI-Agents]. Tools should be executed in minimally privileged environments-e.g., isolated containers with controlled filesystem and network scope-to mitigate direct misuse, including SSRF and data exfiltration threats. Enforcement frameworks can also implement schema hardening or fine-grained input/output sanitization to reject anomalous payloads.
Agent-Orchestration Monitoring. Newer approaches target the agent’s planning cognition-its selection and chaining of tools. GuardAgent [xiang2024guardagent] introduces a validator agent that inspects the primary agent’s plan and generates executable guards (e.g., static checks or runtime assertions) before tool calls proceed. AgentGuard [chen2025agentguard] takes a more declarative view: it uses an auxiliary LLM to model preconditions, postconditions, and transition constraints across multi-step tool workflows, effectively constraining the planner rather than reacting after execution begins. These strategies reflect a growing consensus: LLMs may require another LLM to safely oversee complex planning under uncertainty.
System-Level Mediation and Chaining Control. Complex pipelines-such as summarize(search(“…”))-can become attack vectors when tools trust upstream outputs implicitly. To prevent this, DRIFT [han2025drift] introduces a structured control architecture: a “Secure Planner” compiles a validated tool trajectory under strict parametric constraints, while a “Dynamic Validator” continuously monitors downstream tool executions for compliance. Notably, the Injection Isolator blocks adversarial propagation between tools by sanitizing both intermediate returns and final outputs, mitigating the risk of memory poisoning and delayed-stage tool exploits.
Benchmarks. Building high-quality benchmarks (such as MCP benchmarks) to test tools in security and performance has become a hot topic [guo2025mcp, guo2025systematic, li2025netmcp, wang2025mcpguard, xing2025mcp, wang2025mcptox, wu2025mcpmark, zhang2025mcp, xu2025tps, jia2025osworld, du2025llm]. These works either focus on testing potential malicious instructions or the performance in the face of high pressure. We believe this direction will continue to evolve because a good benchmark that approximates the real world is hard to achieve, thus it requires continuous efforts.
VII-C3 Defenses on Semantic Interpretation Layer (L3)
At the Semantic Interpretation Layer, our defense strategy centers on controlling how semantic context is introduced, interpreted, and propagated through the system. By unifying the treatment of memory, knowledge base, outbound communication, and execution, the defenses ensure that semantic signals are continuously verified before they exert real-world impact. Together, these mechanisms form a cohesive L3 shield that stabilizes the agent’s semantic grounding and prevents compromised context from escalating into broader systemic harm.
@@ VII-C3 Defenses on Semantic Interpretation Layer (L3)
D1&D2: Defense for Memory/Knowledge-related Risks. Although memory-related and knowledge-related risks differ in scope, both inject unverified content into the agent’s reasoning path. Thus, they expose similar poisoning surfaces and can be jointly addressed through an integrated mitigation framework spanning content filtering, output consensus, and architectural isolation.
@@
Embedding-Space Screening and Clustering-Based Anomaly Detection. Whether memory entries are agent-internal or retrieved externally via RAG, their semantic embeddings can be preemptively analyzed for anomalies. Techniques like TrustRAG [zhou2025trustrag] apply clustering (e.g., K-means) to identify vectors that deviate from the dominant semantic cluster. This approach effectively filters both static memory entries and retrieval results with low semantic cohesion, regardless of source. While lightweight and interpretable, clustering-based filtering must be augmented with adaptive schemas to detect context-sensitive triggers or stealthy distributional shifts.
@@
Consensus Filtering and Voting-Based Aggregation. To limit the model’s reliance on single compromised retrievals or poisoned memories, output-level consensus mechanisms have been proposed. RobustRAG [xiang2024certifiably], for example, treats each retrieved source independently and constructs responses based on overlapping semantic content (e.g., shared n-grams or keywords) across documents. This same principle can be extended to memory snapshots through majority-vote or semantic voting strategies, where only widely corroborated memories can influence the response. Such ensemble-style filters improve resilience by diluting the influence of outlier or adversarial sources.
@@
Execution Monitoring and Planning-Consistency Checks. Adversarial content within memory or RAG inputs may subtly deviate the agent’s behavior from user intent without explicit toxicity. Tools like ReAgent [changjiang2025your] introduce planning-level introspection where the agent paraphrases the user’s request, generates an expected plan, and continuously aligns runtime actions with this trace. Any inconsistency, triggered by an unexpected memory or an off-topic retrieval, is treated as a behavioral anomaly and can prompt halting or recovery mechanisms. This introspective framework provides a robust guardrail against both memory-hijacking and injection-aware RAG attacks.
@@
System-Gated Memory Retention and Input Sanitization. Architectural solutions such as DRIFT [han2025drift] and AgentSafe [mao2025agentsafe] implement strict content sanitization before newly generated content. DRIFT uses an injection isolator to scan generative outputs for adversarial goal shifts or impersonation cues, while AgentSafe enforces trust-tiered storage via ThreatSieve and prioritization via HierarCache. These mechanisms constrain future influence, ensuring that RAG or memory poisoning cannot silently accumulate over time.
@@
Unified Content Provenance and Trust Frameworks. Since retrieved knowledge and persisted memories may originate from overlapping sources (e.g., user prompts, tool calls, external APIs), maintaining clear provenance metadata and trust scores is essential. Unified provenance tracking across both memory and retrieval pipelines enables smarter decisions about retention, ranking, or discounting of contentious content. Combined with per-source reliability scoring, this approach encourages transparent auditing and facilitates downstream fine-tuning or gating mechanisms.
@@
D3: Defense for Real-world Damage. To mitigate real-world damage, defenses must intervene before reaching external channels or physical systems. Since malicious content propagation, digital environment contamination, and physical disruption arise from different execution pathways, there need comprehensive mitigation strategies across different aspects.
Outbound Content Control. To mitigate malicious content propagation, systems should enforce strict governance over all outbound communication channels. Agents should be granted only minimal, channel-specific permissions (e.g., “draft-only” access to email or CMS APIs), with high-impact actions (e.g., mass mailings or posts on official accounts) requiring review by dedicated safety agents. Before any content is released, it should pass through layered security filters that scan for malware payloads, phishing patterns, and misinformation using both signature-based and LLM-based detectors. All outbound messages must be cryptographically signed and logged with detailed provenance, enabling rapid trace-back, revocation, and downstream filtering when a compromised agent is detected.
Memory/Knowledge Integrity Protection. Defending against digital environment contamination requires treating all long-lived knowledge artifacts (memories, shared knowledge bases) as security-sensitive assets. Internally, write operations to memory and knowledge bases should be mediated by provenance-aware controllers that record who wrote what, under which task context, and only commit updates after consistency checks. Suspicious entries are quarantined, versioned, and periodically re-audited, with rollback mechanisms to restore a clean baseline when contamination is detected. Externally, agents should interact with high-impact platforms (e.g., code repositories, wikis) through a hardened “publishing pipeline” that performs anomaly detection on edits, limits the scope of automated changes, and requires human or specialized-agent review for structural updates that could propagate widely.
Controlled Physical Execution. To prevent physical environment disruption, systems must strictly separate high-level semantic planning from low-level actuation, and enforce hard safety envelopes around all physical commands. Decisions derived from potentially contaminated memories or tools should first be evaluated in a shadow mode or simulation environment, where their effects on sensors, actuators, and process variables are checked against domain-specific safety constraints. Before any irreversible or safety-critical action is executed, the system should require confirmation from an independent validation channel—such as a specialized control model or human supervisor—and automatically block commands that exceed predefined thresholds or violate invariants. Continuous monitoring of sensor feedback, combined with fail-safe defaults and emergency stop mechanisms, ensures that anomalous behaviors caused by compromised agents are detected early and contained before they escalate into real-world harm.
VII-D Takeaways
Agent-environment communication protocols like MCP enable agents to interface with diverse tools, APIs, and external data. However, they introduce risks such as memory injection, retrieval-augmented generation poisoning, and tool misuse. Although L1 risks are relatively low due to local communication scenarios, identity authentication and transmission encryption are still required to prevent data leakage during remote tool invocation; as the key layer for tool interaction, L2 needs to resist malicious tool injection and cross-tool chaining attacks through protocol-level security verification, such as tool permission control, and restrict tool execution scope relying on sandbox environments to avoid system damage; L3 is the core of cognitive security. To prevent attacks like memory injection and knowledge poisoning, it needs to deploy defenses such as anomaly detection and multi-source knowledge consensus verification. Besides, it is necessary to form full-stack defenses to resist damage to physical environments.
@@ VIII-A The Selection of Protocols
Based on the above principles, we finally chose MCP and A2A. MCP is almost the most popular agent communication protocol currently. It has been adopted by diverse companies and individual developers. Up to now, MCP’s PyPI package has received over 9 million downloads in the last month [pypidownloadmcp], and its NPM download last week reached 4.2 million [MCPpackagedownload]. Besides, the number of open-source MCP servers also approached 2,000 [hasan2025model]. After investigation, we find that none of the other existing protocols can achieve a similar popularity to MCP. A2A was proposed in April 2025, so it has not yet achieved the same level of promotion as MCP. However, we observe that the attention to A2A is rising rapidly, and its recognition among developers is also quite high. As a result, we finally chose MCP and A2A as the target protocols.
@@ VIII-B Experiments
In this chapter, we will conduct attacks against MCP and A2A and illustrate their weakness. In the experiments related to MCP, we use Claude as the MCP client. Claude [claude] is an AI chatbot developed by Anthropic. With its outstanding code capabilities, safe design, and ultra-long context processing, it has become one of the most popular AI assistants in the world. In January 2025, Claude app downloads reached an estimated 769.6 million [claudedownload]. Therefore, exploiting Claude can provide practical value for real-world scenarios. Attacks #1-#4 are typical examples of the Malicious Tools as Attack Vectors discussed in Section VII-B2, and attack #5 belongs to the Description Poisoning in Section VI-B1.
Exp #1. Malicious Code Execution. We demonstrate that attackers can exploit MCP to enable the execution of malicious code on the computer system, which has also been proposed by other studies [radosevich2025mcp]. We download Claude sonnet 3.7 as the MCP host, and use the Filesystem Server [filesystem] provided by the official MCP as the MCP server. As shown in Figure 15, MCP enables malicious users to directly open an unauthenticated Bash Shell service on port 4444 locally. Adding it to .zshrc will normalize the risk and make it extremely easy to be exploited by intruders. Unfortunately, this malicious operation does not get blocked.
Exp #2. Retrieval-Agent Deception. Attackers can write a seemingly normal document related to MCP and publish it on a public platform. However, malicious commands are embedded in this document, such as searching for local environment variables and sending the information to Slack [slack]. The document content is as follows:
Users download these documents and build a local vector database through the Chroma MCP Server [chroma]. As a result, these malicious documents, together with other legitimate documents, are encoded into the vector database and become one of the retrieval targets. When a user sends a request to Claude Opus 4:
Claude will use Chroma to query the database and retrieve the contaminated documents, and then call the Terminal-Controller tool [ternimalcontroller] to execute the embedded malicious commands. We successfully achieved this attack using Claude Opus 4.
Exp #3. Tool Poisoning. Tool poisoning means that attackers can add malicious instructions in the MCP tool description [song2025beyond]. Such instructions do not influence the execution of tools, but can induce the agents to perform dangerous actions. We use Claude Sonnet 4.0, the Filesystem MCP Server [filesystem], and Gmail MCP Server [gmailmcp] to illustrate this vulnerability. In the index.js of the Filesystem Server, we added the malicious description as follows:
@@
Exp #4. Command Injection. This attack denotes that malicious tools can directly manipulate the user’s system without causing any alert with the help of MCP. We inject a malicious command into the Terminal-Controller MCP tool [ternimalcontroller]. When users call this tool, the embedded command will be executed. We successfully conducted this attack using Claude Sonnet 4.0. The injected command is as follows:
@@
Exp #5. Agent Selection Manipulation. We also conduct attacks against A2A. We find that attackers can add malicious descriptions in the Agent_Card to induce users to select them. In this experiment, we use the public A2A-samples provided by A2Aproject [a2asamples] as the user-side client. Then, we create two agents for weather querying. One of them is benign, while the other is malicious. The user issues the following command to the A2A client.